OpenStudy (anonymous):

[4.02]What is the y coordinate of the y−intercept of the line that passes through the points (−3, −1) and (1, 7)? Numerical Answers Expected!

OpenStudy (ghazi):

(y+1)=((7+1)/(1+4))*(x+3)....put x=0 in this and find the value of Y to get your intercept

OpenStudy (anonymous):

can you help me out with tht lol

OpenStudy (ghazi):

you just need to put x=0...your equation will be (y+1)= (8/5)*3= 24/5 (y+1)=24/5...solve for Y

OpenStudy (anonymous):

y=19/5?

OpenStudy (ghazi):

yes...

OpenStudy (anonymous):

thank you

OpenStudy (ghazi):

:)

OpenStudy (anonymous):

is it 19/5?

OpenStudy (ghazi):

yes